# Testing stimulated emission photon directions

### v4

@@ Abstract
While naively laser only causes excitation of external target, for example STED microscopy or backward ASE in optical amplifiers, demonstrate it can also cause its deexitation, however, under uncommon condition of being prepared as excited. These two causalities are governed by absorption-stimulated emission pair of equations, and swap places in perspective of CPT symmetry, however, it means photon direction of stimulated emission should be opposite to usually assumed. This article discusses various arguments and proposes simple direct tests to experimentally verify existence of such backward photon trajectories, complementing consequent forward textbook trajectories. There are already indirect experimental arguments, like backward ASE, lack of Rabi cycle for free electron laser, or recent ”negative time evidence” experiment in the news. Depending on the result: it should lead to new applications like 2WQC more symmetric quantum computers, or macroscopic violation of CPT symmetry, so far probably tested only in microscopic settings. Also proposed forward/backward/bidirectional photon sources are discussed.
@@ I Introduction
Figure 1: Top: recent widely commented experiment by Steinberg group [1] has observed medium response (measured phase of 85Rb) before and after the applied impulse, proposing ”photon can spend a negative amount of time in an atom cloud” type of explanation, which does not seem physical. In contrast, as discussed here, this might be a result of just stimulated emission, required as some of the atoms were excited by magneto-optical trap (MOT) lasers, and if being CPT analog of absorption, should have opposite delay sign.
@@
Figure 2: Top: CPT symmetry says that if there exists ”laser causes excitation” scenario, also ”laser causes deexciation” should exist. Being CPT symmetry analogs, they should have opposite photon direction and delay sign - we would like to verify it experimentally and apply if successful. For example in STED microscopy diode lasers cause both. Bottom: proposed STED-like test of this photon direction(s): continuous laser excites fluorescent dye being approximately 3-level medium, and impulse laser causes its deexcitation through stimulated emission - its photon direction can be found from $\Delta t=(d\pm l)/c$ delay between laser impulse and observed reduced intensity by fast photodetector focused on dye spontaneous emission. Textbooks assumption would mean ${}^{\prime}+^{\prime}$ sign here, while CPT symmetry requires ${}^{\prime}-^{\prime}$ sign instead, hence this is also macroscopic test of CPT symmetry, complementing dozes of earlier microscopic tests [3].
Figure 3: In commercially available e.g. erbium-doped optical fiber amplifiers (EDFA or alternatives) [4] there is observed strong backward ASE (amplified spontaneous emission also called superluminescence) - increased emission from pumped doped fiber toward laser, exactly as in the sought CPT analog of ”laser causes excitation” effect. Analogously to STED-like test, we can verify (and apply if successful) this photon direction by measuring delay between the impulse and reduced population of excited atoms of external target, e.g. by monitoring its spontaneous emission. For this purpose, the shown standard forward optical isolator should be removed, or better rotated to backward position. As erbium has very long millisecond-scale lifetimes (should be shortened for practical applications), there might be required cascade photodiode.
@@ II-B Alternative STED-like tests
We could also test suggestions of better transparency for backward beam, e.g. by putting (non-excited) obstacle on the way to depletion laser in STED-like setting. It also brings hope for effects on excited targets already for shielded synchrotron sources. Even if such photons could not cross the obstacle, this effect could be used e.g. for information transfer by trying to detect such additional directional photons from excited target behind obstacle, or monitoring its population level.
@@ III-A Bidirectional and forward/backward beams
Figure 6: Top: ring laser uses optical isolator to enforce unidirectional photon circulation, which would be reversed in CPT perspective - switching the action of absorption and stimulated emission equations, suggesting only one of them should act on external targets - let us refer to them as forward/backward beam. Analogously for just adding optical isolator, or synchrotron sources like wiggler, undulator, free electron laser - reversing electron and photon trajectory in CPT perspective. Center: STED microscopy (image source) showing that, at least for diode lasers, both equations act on external target, hence we can call it bidirectional beam. We usually focus on absorption by external target, because stimulated emission requires its initial excitation, while without such preparation usually $N_{2}/N\approx 0$. In Rabi cycle laser periodically causes excitation and deexitation of target, suggesting requirement of bidirectional beam. Its experimental support is lack of Rabi oscillations for free electron laser [14] - leading to $N_{2}\approx N$ nearly complete excitation in Autler-Townes effect (dynamic Stark), instead of $N_{2}\approx N/2$ for damped Rabi cycle. Bottom: while CPT symmetry of physics is required by local Lorentz-invariant QFT, e.g. switching forward/backward beams, in theory general relativity allows its realization by travelling through nonorientable wormhole [15], bringing tangible thought experiment valuable for intuitions.
@@
As noticed in Fig. 4, forward/backward beams might have completely different transparency as they act only on $N_{1}$ deexcited / $N_{2}$ excited atoms. In other words, while textbook view on stimulated emission means amplification of forward beam by going through excited medium, its CPT analog is amplification of backward beam by going through deexcited medium (as in this Figure). Therefore, as everyday matter usually has $N_{2}/N\approx 0$, backward beam might have much better penetration, allowing for novel applications.
@@
III-B Experimental consequences like Rabi cycle
To test the above with STED-like setting, as mentioned we can replace depletion laser with forward/backward beam e.g. by optical isolator - only backward should lead to depletion. If stimulated emission-absorption use opposite photon trajectories, we could also try to split them exploiting delay with fast shutter or switch, or maybe exploiting transparency absorbing only forward beam.
Another type of consequence is $N_{1},N_{2}$ population enforced by laser in its external target. For (symmetric) bidirectional beam acting with both equations, we could at most reach 1/21/$2-1/$ 2 - 1 / 2 population, and it is indeed well known as the final population of damped Rabi cycle [18]: using periodically stimulated emission and absorption equations. In Fig. 7 there is suggested hydrodynamic analog of Rabi cycle (another can be found in [19]), which should exist through EM-hydrodynamics analogy as written, like the shown resonator coupled with external object exchanging momentum in both directions, what might be also appropriate intuition for bidirectional lasers, using cyclically positive and negative radiation pressure [20]. Rabi cycle in CPT perspective should be practically the same - bringing another argument for photons travelling there only between laser and target.
However, using only forward beam, we should be able to reach $N_{2}\approx N$ nearly complete excitation, and such effect has turned out to be known in literature: despite many attempts and theoretical prediction, they were not able to obtain Rabi oscillation with free electron laser [14] - instead getting Autler-Townes effect (dynamics Stark) leading to $N_{2}\approx N$, instead of $N_{2}\approx N/2$ for damped Rabi oscillations. There should be also possible asymmetric bidirectional beams e.g. by adding multiple beams, or applying imperfect optical isolators - with different powers of forward/backward beams, for example leading to damped Rabi cycle approaching a different population.
@@ III-B Experimental consequences like Rabi cycle
Figure 7: Top: Mathematical analogy between electrodynamics and hydrodynamics (superfluid - no viscosity) from [21]. Center: therefore, Rabi cycle also needs hydrodynamical analog, tank of liquid with standing wave seems analogous to laser resonator. Being semi-transparent/semi-open in one direction, it should cause left-right swimming of a probe particle outside, through exchange of momentum acting in both directions. In CPT perspective it looks the same, also with momentum exchange in both directions. Bottom: while the above bidirectional source is resonator kind of extending to external target e.g. for Rabi oscillations, there are also unidirectional sources. In this analogy, synchrotron radiation is kind of water stream allowing only unidirectional momentum transfer. Optical isolator analog could be some unidirectional valve, analog of ring laser a pump enforcing unidirectional circulation - with reversed circulation in CPT perspective. Radiation pressure is a vector: can be also negative [20].
III-C Some possible new applications
@@
If confirmed, the above considerations could lead to many new applications of e.g. standard optical fiber equipment, especially for the currently neglected backward beam, like:
@@
  • Forward beam (e.g. by adding optical isolator) could allow for state preparation of quantum computers to nearly completely excited $N_{2}\approx N$ as in [14], backward beam for its CPT analog (postparation) for 2WQC: two-way quantum computers [7] in theory allowing to solve NP problems [22].
@@
  • If we could send information this way, there could be e.g. built time-loop computers (Section V of [6]) for example using backward beam to send chip output back to its original input with negative time delay, also in theory allowing to solve NP problems. Connecting such multiple negative delay transmitters e.g. into a cycle, could allow to also reach macroscopic negative time differences.
  • Backward camera e.g. exploiting transparency or for astronomical observations - focused on target’s backward beam, while standard camera is focused on its forward beam. It would replace CMOS/CCD matrix behind the lens, with matrix/surface of continuously excited approximately 3-level medium like dye, plus some monitoring of its population map - e.g. through spontaneous emission, or EIT (electromagnetically induced transparency).
@@
  • Backward beam might be transparent for $N_{2}=0$ medium, what could allow for much better penetration than forward (transparent for $N_{1}=0$) - e.g. for backward camera, various scanners, communication through obstacles.
@@
  • Backward beam to stimulate specific chemical processes: releasing photons of chosen energy/polarization, e.g. for chemical/technological processes, energy production, or medical therapies - e.g. locally reduce lifetime of ATP molecule (into thermal energy also heating), or inhibit biochemical processes by causing deexcitation of crucial molecules, for example with multiple beams intersecting in cancer tissue, especially if better transparency is confirmed.
  • Backward beam for 3D printing/reversed lithography: excite material and precisely stimulate its deexcitation - controlling position and time, also frequency to choose one of multiple possible transitions e.g. some high energy transition to locally reduce temperature, maybe polarization.
@@
IV Conclusions and further work

### v6

@@ Abstract
While naively laser only causes excitation of external target, e.g. Rabi cycle, STED microscopy or backward ASE in optical amplifiers demonstrate it can also cause its deexitation, however, under uncommon condition of being prepared as excited. These two causalities are governed by absorption-stimulated emission pair of equations, and swap places in perspective of CPT symmetry, however, it means photon direction of stimulated emission should be opposite to usually assumed. This article discusses various arguments and proposes simple direct tests to experimentally verify existence of such backward photon trajectories, complementing consequent forward textbook trajectories. There are already indirect experimental arguments, like backward ASE, lack of Rabi cycle for free electron laser, or recent ”negative time evidence” experiment in the news. Depending on the result: it should lead to new applications like 2WQC more symmetric quantum computers, or macroscopic violation of CPT symmetry, so far probably tested only in microscopic settings. Also proposed forward/backward/bidirectional photon sources are discussed.
@@ I Introduction
Figure 1: Top: recent widely commented experiment by Steinberg group [1] has observed medium response (measured phase of 85Rb) before and after the applied impulse, proposing ”photon can spend a negative amount of time in an atom cloud” type of explanation, which does not seem physical.
Bottom: instead, as discussed here, it might be a result of just stimulated emission due to EM coupling of laser resonator and atoms as tiny resonators, required as some of the atoms were excited by magneto-optical trap (MOT) lasers, and if being CPT analog of absorption, should have opposite delay sign.
@@
Figure 2: Top: CPT symmetry says that if there exists ”laser causes excitation” scenario, also ”laser causes deexciation” should exist. Being CPT symmetry analogs, they should have opposite photon direction and delay sign - we would like to verify it experimentally and apply if successful. For example in STED microscopy diode lasers cause both.
Bottom: proposed STED-like test of this photon direction(s): continuous laser excites fluorescent dye being approximately 3-level medium, and impulse laser causes its deexcitation through stimulated emission - its photon direction can be found from $\Delta t=(d\pm l)/c$ delay between laser impulse and observed reduced intensity by fast photodetector focused on dye spontaneous emission. Textbooks assumption would mean ${}^{\prime}+^{\prime}$ sign here, while CPT symmetry requires ${}^{\prime}-^{\prime}$ sign instead, hence this is also macroscopic test of CPT symmetry, complementing dozes of earlier microscopic tests [3].
Figure 3: In commercially available e.g. erbium-doped optical fiber amplifiers (EDFA, or simpler semiconductor SOA, etc.) [4] there is observed strong backward ASE (amplified spontaneous emission) - increased emission from pumped doped fiber toward laser, exactly as in the sought CPT analog of ”laser causes excitation” effect. Top:Analogously to STED-like test, we can verify this photon direction by measuring delay between the impulse and reduced population of excited atoms of external target, e.g. by monitoring its spontaneous emission. For this purpose, the shown standard forward optical isolator should be removed, or better rotated to backward position. As erbium has very long millisecond-scale lifetimes, there is required fast photodiode. Fig. 7 contains further simplified test. Bottom: simplified universal setting as in [5], more practical to measure backward ASE delay, separation, transparency, and for their potential applications.
@@
Center: basic argument for ()($\leftarrow$)( ← ) is CPT symmetry of physics: transforming ”laser causes target excitation” scenario into ”CPT(laser) causes CPT(target) deexcitation”, also switching the stimulated emission-absorption equations, governed by the same (CPT symmetric) Einstein’s coefficients $B_{12}=B_{21}$.
Bottom: schematic explanation that CPT view ()($\leftarrow$)( ← ) also allows to conclude textbook ()($\rightarrow$)( → ) picture, hence both can be true. For photons travelling in ”laser causes excitation/deexcitation” CPT analogous scenarios, we add a medium on the way: looking at transparency in the latter, stimulated emission should fully act on the target, but some photons it produces should be absorbed by medium - returning from CPT perspective it becomes textbook amplification by going through excited medium. All should be coupled by standing waves as in Fig. 7, stimulating/amplifying all these photon emissions.
@@ II-B Alternative STED-like tests
We could also test suggestions of better transparency for backward beam, e.g. by placing (non-excited) obstacle on the way to depletion laser in STED-like setting. It also brings hope for effects on excited targets already for shielded synchrotron sources. Even if such photons could not cross the obstacle, this effect could be used e.g. for information transfer by trying to detect such additional directional photons from excited target behind obstacle, or monitoring its population level.
@@
Figure 6: Top: ring laser uses optical isolator to enforce unidirectional photon circulation, which would be reversed in CPT perspective - switching the action of absorption and stimulated emission equations, suggesting only one of them should act on external targets - let us refer to them as forward/backward beam. Analogously for just adding optical isolator, or synchrotron sources like wiggler, undulator, free electron laser - reversing electron and photon trajectory in CPT perspective.
Center: STED microscopy (image source) showing that, at least for diode lasers, both equations act on external target, hence we can call it bidirectional beam. We usually focus on absorption by external target, because stimulated emission requires its initial excitation, while without such preparation usually $N_{2}/N\approx 0$. In Rabi cycle laser periodically causes excitation and deexitation of target, suggesting requirement of bidirectional beam. Its experimental support is lack of Rabi oscillations for free electron laser [15] - leading to $N_{2}\approx N$ nearly complete excitation in Autler-Townes effect (dynamic Stark), instead of $N_{2}\approx N/2$ for damped Rabi cycle.
Bottom: while CPT symmetry of physics is required by local Lorentz-invariant QFT, e.g. switching forward/backward beams, in theory general relativity allows its realization by travelling through nonorientable wormhole [16], bringing tangible thought experiment valuable for intuitions.
@@ III-A Bidirectional and forward/backward beams
As noticed in Fig. 4, forward/backward beams might have completely different transparency as they act only on $N_{1}$ deexcited / $N_{2}$ excited atoms. In other words, while textbook view on stimulated emission means amplification of forward beam by going through excited medium, its CPT analog is amplification of backward beam by going through deexcited medium (as in this Figure). Therefore, as everyday matter usually has $N_{2}/N\approx 0$, backward beam might have much better penetration, allowing for novel applications. However, coupling standing wave might need to reach the target, what could essentially weaken such effect.
@@
III-B Optical cavity and delocalized standing EM wave coupling
In vacuum Rabi cycle [19] e.g. a single atom periodically exchanges energy with optical cavity - cycling energy between atomic orbital, and standing EM wave of cavity, with energy density more localized in the former case. This difference is similar to between laser and radio communication - the former uses localized optical photons travelling in one direction, the latter delocalized e.g. spherical EM waves travelling in all directions.
While both are usually described with Fock space, this is abstract effective description of pertubative approximation. Like ”apple + apple = 2 apples” correct algebra, for deeper dependencies requiring to ask about structure, field configuration in non-perturbative picture, here mainly of electromagnetic fields - again with higher energy density when localized in atom e.g. during vacuum Rabi oscillations.
Being different types of EM field configurations, their interactions might also differ, for example with sophisticated devices like optical isolator - specialized to block localized optical photons travelling in one direction, but not necessarily delocalized standing EM waves for coupling - what should be tested, e.g. with setting as in Fig. 7. Finding material which acts as a mirror for optical photons, but is semi-transparent for delocalized EM waves for coupling, would be perfect to close resonator of potential backward lasers: only amplifying target deexcitation, e.g. for radiotherapy as in Fig. 10.
III-C Rabi cycle as coupling, its asymmetrization
Generally, in Rabi cycle as in Fig. 7, analogously to coupled pendula (e.g. [20, 21, 22, 23]), there is observed coupling of two resonators periodically exchanging energy and momentum in both directions - in practice using mainly lasers as pumped directional resonators, atoms as tiny isotropic resonators, or cavities leading to standing waves/eigenmodes.
As electromagnetism and hydrodynamics are governed by nearly the same equations, it is also valuable to think about Rabi cycle from hydrodynamics perspective as in Fig. 8: laser as standing wave extending outside resonator due to semi-transparent wall, exchanging energy and momentum with external objects in both directions - with positive and negative pressure, both exist also for radiation [24].
Such bidirectional energy exchange for e.g. laser-atoms coupling is mainly realized by optical photons, hence we can view standard e.g. diode laser as bidirectional beam - causing both excitation of target, but also deexcitation - e.g. in Rabi cycle, STED microscopy, or backward ASE of EDFA.
Being able to introduce asymmetry in this coupling could lead to new applications. For example free electron laser (FEL) does not have such resonator coupling, hence instead of damped Rabi [25] leading to $N_{2}\to N/2$, FEL experiments [15] show $N_{2}\to N$ only excitation with unidirectional: forward beam - photon exchange only in one direction, acting only with absorption equation on the target. CPT symmetry suggests that looking at such synchrotron radiation beam from CPT perspective, photon beam created by reversed electron trajectory should act as backward beam - causing only target deexctiation, acting only with stimulated emission equation.
As synchrotron radiation sources are large and costly, for practical reasons we should search ways to separate absorption and stimulated emission equations with different tools like optical isolator, ring laser - to be tested.
Figure 7: Top:coupled pendula (image source) as analog of Rabi cycle (e.g. [20, 21, 22, 23]) - to view it as coupling of e.g. laser as pumped directional resonator, and atom as tiny isotropic resonator, or standing waves of optical cavity as resonator e.g. through Jaynes-Cummings model [26].
Center: vacuum Rabi cycle [19] - observed oscillations between optical cavity resonator, and atom as tiny resonator - of energy between being localized in atom (high energy density), and delocalized standing EM wave (low energy density) - having different EM fields, like localized laser vs delocalized radio communication. Laser can enforce Rabi oscillations, periodically causing target excitation and deexcitation - allowing to view these causalities as results of coupling of two resonators, in contrast e.g. to free electron laser without resonator - hence only observing target excitations there [15]: $(N_{2}\to N)$ instead of damped Rabi $(N_{2}\to N/2)$. Being able to add some asymmetry to such energy transfer caused by coupling, as discussed further can lead to many applications - e.g. optical isolator brings hope for that: designed to block localized optical photons in one direction, what not necessarily has to be true for delocalized standing wave for coupling - having lower energy density.
Bottom: proposed test if optical isolator blocks this EM coupling - monitoring backward ASE: amplified emission from erbium toward laser, for 3 optical isolator cases. It is observed without optical isolator (’none’ case), backward ASE photons are often blocked by $\rightarrow$ forward placed optical isolator. The main question is if $\leftarrow$ backward placed optical isolator leaves backward ASE - suggested both by CPT symmetry (switching photon and optical isolator direction), and coupling through delocalized EM wave.
Figure 8: Top: Mathematical analogy between electrodynamics and hydrodynamics (superfluid - no viscosity) from [27].
@@ III-C Rabi cycle as coupling, its asymmetrization
Center: therefore, Rabi cycle also needs hydrodynamical analog, tank of liquid with standing wave seems analogous to laser resonator. Being semi-transparent/semi-open in one direction (also in Fig. 7), it should cause left-right swimming of a probe particle outside, through exchange of momentum acting in both directions, using positive/negative pressure. In CPT perspective it looks the same, also with energy/momentum exchange in both directions.
Bottom: while the above bidirectional source is resonator extending to external targets, coupling with them through standing wave e.g. for Rabi oscillations, there are also unidirectional sources. In this analogy, synchrotron radiation is kind of water stream allowing only unidirectional momentum transfer. Optical isolator analog could be some unidirectional valve, analog of ring laser: a pump enforcing unidirectional circulation - with reversed circulation in CPT perspective. Radiation pressure is a vector: can be also negative [24], suggesting to build more symmetric and powerful two-way microfluidic, photonic computers as in Fig. 9.
IV Some potential new applications
@@
If confirmed, the above considerations could lead to many new applications, especially for the currently neglected backward beam - optimistically of standard e.g. optical fiber equipment (resonator coupling) or ring lasers, pessimistically of synchrotron radiation (forward/backward sources), like:
@@
  • Forward beam, e.g. from synchrotron radiation, could allow for state preparation of quantum computers to nearly completely excited $N_{2}\approx N$ as in [15], backward beam for its CPT analog (postparation) for 2WQC [8] as in Fig. 9 in theory allowing for up to exponential speedup and better stability ([28, 29]).
@@
  • If we could send information this way, there could be e.g. built time-loop computers (Section V of [7]) for example using backward beam to send output of chip back to its original input with negative time delay, also in theory allowing to solve NP problems. Connecting such multiple negative delay transmitters e.g. into a cycle, could allow to also reach macroscopic negative time differences.
@@
  • Backward beam might be transparent for $N_{2}=0$ medium, what could allow for much better penetration than forward (transparent for $N_{1}=0$) - e.g. for backward camera as in Fig. 11, various scanners, communication through obstacles. Finding material being mirror for optical photons, but semi-transparent for standing EM waves for coupling, it would be perfect to close resonator of potential backward laser.
  • Backward beam to stimulate specific chemical processes: releasing photons of chosen energy (and polarization), e.g. for chemical/technological processes, energy production, or medical therapies as in Fig. 10 - e.g. locally reduce lifetime of ATP molecule (stimulate decay also releasing this energy locally heating), or inhibit biochemical processes by amplifying deexcitation of crucial molecules, for example with multiple beams intersecting in cancer tissue, especially if better transparency is confirmed.
  • Backward camera maybe exploiting transparency and/or opposite delay e.g. for astronomical observations - focused on target’s backward beam, while standard camera is focused on its forward beam. It would replace CMOS/CCD matrix behind the lens, with matrix/surface of continuously excited approximately 3-level medium like dye, plus some monitoring of its population map - e.g. through spontaneous emission, or EIT (electromagnetically induced transparency).
@@
  • Backward beam for 3D printing/reversed lithography: excite material and precisely stimulate deexcitation - controlling position, time, polarization, also frequency to choose one of multiple possible transitions e.g. of high energy to locally reduce temperature (cooling), of energy regained thermally.
@@
Figure 9: Original motivation: (example of) two-way quantum computer (2WQC) [8] as adding a process which in CPT perspective becomes the original state preparation process. For photonic QC state preparation is usually performed by laser impulse, suggesting to just add ’reversed’ impulse: delayed by propagation time, and behind backward placed optical isolator - this way from CPT perspective photon trajectories should be reversed. In comparison to standard 1WQC, in theory 2WQC allows for up to exponential speedup e.g. to solve NP problems, also better stability and error correction [28, 29]. Such classical 2WCC photonic computer on continuous beam might be also interesting.
Figure 10: Being able to split absorption and stimulated emission equations, for example with CPT analog of synchrotron radiation: beam resulting from reversed electron trajectories as in Fig. 6, or maybe standard laser with backward placed optical isolator as in Fig. 7, one of many new applications could be radiotherapy trying to locally amplify emission of photons of chosen energy and polarization. For example to try to locally increase degradation of energy carrying molecules like ATP (if having emission spectrum) - to starve cancer tissue, and heat up by releasing this energy. Intersecting multiple such beams, maybe additionally rotating them for fixed intersection point, we could make the effect there orders of magnitude stronger then for intermediate tissues, to make it negligible outside of targeted tissue. Different issues could be also targeted this way, e.g. trying to disable toxic molecules, inhibit chosen biochemical pathways by stimulating deexctiation of crucial molecules e.g. of pathogens, or just for local cooling, etc.
Figure 11: Top: Backward camera observing backward ASE caused by the target, especially interesting if confirmed better transparency and/or opposite delay sign. Can be e.g. based on (multi-mode) fiber with circulators as in [5] and Fig. 3, or more practical with semitransparent mirror and CCD/CMOS - both need some continuously pumped medium (maybe from some display technology). Alternatively one could monitor spontaneous emission of the pumped medium.
Bottom: Backward lighted camera: applying backward ASE effect on the target to amplify its emission in chosen spectrum toward camera (e.g. thermal also cooling), or on laser medium used to send quantum cryptography protocols like BB84, to increase its number of produced photons and steal/read some.
V Conclusions and further work